$SPAD/src/input f07adf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
m:=4
n:=4
lda:=4
a:Matrix SF:=
 [[1.8 ,2.88 ,2.05 ,-0.89 ],_
  [5.25 ,-2.95 ,-0.95 ,-3.8 ],_
  [1.58 ,-2.69 ,-2.9 ,-1.04 ],_
  [-1.11 ,-0.66 ,-0.59 ,0.8 ]]
result:=f07adf(m,n,lda,a)

———-

References^{[1] nothing}

4